# From Stochasticity to Signal: A Bayesian Latent State Model for Reliable Measurement with LLMs

### v1

@@ 5.3 Model Comparison
Shown in Table 2, the extended model with ten LLM ratings for each call outperforms the rest of the methods in terms of the correlation with the latent dissatisfaction status (0.803). When the heterogeneity in error rates is captured, the extended model with five rounds of LLM outputs outperforms the base model with ten rounds of outputs. Both extended and base methods have improved accuracies in estimating the true latent dissatisfaction as the input rounds increases. The majority vote approach with all ten rounds of input merely outperforms the proposed methods with one single round of output.
When the impact parameters are of interest, Table 3 shows that the performance of the majority vote approach further deteriorates by providing the least accurate estimates. The extended method produces more accurate estimates than the base method conditioning on all three choices of output rounds. However, the accuracy in recovering the impact parameters does not increase along with the output rounds. The extended method with five rounds of LLM output provides the best estimate of both impact parameters across all tested model settings.
This phenomenon illustrates a critical tradeoff in mixture models with dual purposes of inference. The likelihood is a mixture of two components: a latent outcome regression component (the logistic regression modeling the true state, where $\tau$ and $\eta$ are estimated) and a measurement error component (the binomial likelihood of the observed ratings, conditional on the true state). As the number of LLM output rounds increases, the binomial likelihood becomes substantially more concentrated. Consequently, the model’s objective becomes dominated by fitting the highly flexible individualized error rate functions to the precise $k_{c}/N_{c}$ ratios observed for each unit. This can lead to a form of parameter confounding, where variance that should be attributed to the treatment effect is instead absorbed by the measurement component. The model mis-attributes a portion of the true causal effect to nuanced patterns in measurement error, resulting in biased estimates for $\tau$ and $\eta$. The N=5 scenario appears to strike an optimal balance, where the measurement component is informative but does not completely overwhelm the latent outcome regression component of the likelihood.
@@
Heterogeneous (N=10) 0.803
Homogeneous (N=10) 0.702
Heterogeneous (N=5) 0.723
Homogeneous (N=5) 0.670
Heterogeneous (N=1) 0.485
Homogeneous (N=1) 0.455
Majority Vote (N=10) 0.662
@@
True Value -0.80 -0.086
Heterogeneous (N=10) -0.51 -0.059
Homogeneous (N=10) -0.31 -0.054
Heterogeneous (N=5) -0.69 -0.072
Homogeneous (N=5) -0.48 -0.070
Heterogeneous (N=1) -0.52 -0.050
Homogeneous (N=1) -0.43 -0.049
Majority Vote (N = 10) -0.29 -0.041

### v3

@@ 5.3 Model Comparison
As shown in Table 2, the extended model with ten LLM ratings for each call outperforms the rest of the methods in terms of the correlation with the latent dissatisfaction status (0.798). When the heterogeneity in error rates is captured, the extended model consistently outperforms the base model across different settings of input rounds. Both extended and base methods have improved accuracies in estimating the true latent dissatisfaction as the input rounds increase. The majority vote approach with all ten rounds of input exhibits a comparable performance to the proposed methods with five rounds of output.
When the impact parameters are of interest, Table 3 shows that the performance of the majority vote approach further deteriorates, providing less accurate estimates than our proposed methods with five rounds of output. The extended method produces overall more accurate estimates than the base method conditioning on all three choices of output rounds. However, the accuracy in recovering the impact parameters does not necessarily increase along with the output rounds. For example, both the base method and the extended method with five rounds of LLM output outperform their counterparts with ten rounds of LLM output in terms of the recovery $\tau$.
This phenomenon illustrates a critical tradeoff in mixture models with dual purposes of inference. The likelihood is a mixture of two components: a latent outcome regression component (the logistic regression modeling the true state, where $\tau$ and $\eta$ are estimated) and a measurement error component (the binomial likelihood of the observed ratings, conditional on the true state). As the number of LLM output rounds increases, the binomial likelihood becomes substantially more concentrated. Consequently, the model’s objective becomes more dominated by fitting the highly flexible individualized error rate functions to the precise $k_{c}/N_{c}$ ratios observed for each unit. This can lead to a form of parameter confounding, where variance that should be attributed to the treatment effect is instead absorbed by the measurement component. The model misattributes a portion of the true causal effect to nuanced patterns in measurement error, resulting in more biased estimates for $\tau$ and $\eta$. In this case, a ‘medium’ number of LLM rounds might be preferred with the dual purpose of estimating the latent states and impact parameters.
@@
Heterogeneous (N=10) 0.798
Homogeneous (N=10) 0.725
Heterogeneous (N=5) 0.699
Homogeneous (N=5) 0.627
Heterogeneous (N=1) 0.434
Homogeneous (N=1) 0.394
Majority Vote (N=10) 0.659
@@
True Value -0.80 -0.085
Heterogeneous (N=10) -0.86 -0.087
Homogeneous (N=10) -0.61 -0.092
Heterogeneous (N=5) -0.84 -0.077
Homogeneous (N=5) -0.64 -0.086
Heterogeneous (N=1) -0.71 -0.051
Homogeneous (N=1) -0.48 -0.035
Majority Vote (N = 10) -0.52 -0.065